what is the gravational force and columbic force in newton units if you have two electrons in 0.2 angstrom
Answer #1

Here,

distance , d = 0.20 A = 2 *10^-11 m

gravitational force = G * m^2/d^2

gravitational force = 6.673 *10^-11 * (9.11 *10^-31)^2/(2 *10^-11)^2

gravitational force = 1.385 *10^-49 N

the gravitational force between the electrons is 1.385 *10-49 N

Now, for the coloumbic force

coloumbic force = k * q^2/d^2

coloumbic force = 9 *10^9 * (1.602 *10^-19)^2/(2 *10^-11)^2

coloumbic force = 5.77 *10^-7 N

the coloumbic force between the electrons is 5.77 *10-7 N
Answer #2

The gravitational force and the Coulombic force are two different types of forces that act between objects. The gravitational force is the force of attraction between any two objects with mass, while the Coulombic force is the force of attraction or repulsion between charged particles.

In the case of two electrons separated by a distance of 0.2 angstroms (1 angstrom = 10^-10 meters), the gravitational force between them can be considered negligible because the force of gravity is much weaker than the electromagnetic force. Therefore, we will only calculate the Coulombic force between them.

The Coulombic force between two point charges q1 and q2 separated by a distance r is given by Coulomb's law:

F = (k * q1 * q2) / r^2

where k is Coulomb's constant, which has a value of approximately 8.99 x 10^9 N·m^2/C^2.

For two electrons, the charge of each electron is -1.602 x 10^-19 C. Therefore, the Coulombic force between two electrons separated by a distance of 0.2 angstroms is:

F = (8.99 x 10^9 N·m^2/C^2) * (-1.602 x 10^-19 C)^2 / (0.2 x 10^-10 m)^2 F ≈ 2.31 x 10^-28 N

Therefore, the Coulombic force between two electrons separated by a distance of 0.2 angstroms is approximately 2.31 x 10^-28 N.


answered by: Hydra Master